# The Inertial Spin Model of flocking with position-dependent forces

### v2

@@ Abstract
We consider the inertial spin model (ISM) of flocking and swarming. The model has been introduced to explain certain dynamic features of swarming (second sound, a lower than expected dynamic critical exponent) while preserving the mechanism for onset of order provided by the Vicsek model. The ISM has only been formulated with an imitation (“ferromagnetic”) interaction between velocities. Here we show how to add position-dependent forces in the model, which allows to consider effects such as cohesion, excluded volume, confinement and perturbation with external position-dependent field. We study numerically a one-particle case.
@@ I Introduction
Although the Vicsek model is successful in explaining the thermodynamics of flocking, it is not suitable to interpret certain dynamical features related to the presence of inertial effects. The finding of wave-like propagation of direction information during turns in starling flocks Attanasi et al. (2014a) led to the proposal of the inertial spin model (ISM) Cavagna et al. (2014), which we consider here. The model is described in detail in Sec. II.1, but it is essentially the Vicsek model endowed with a Hamiltonian-like (second order) dynamics, so that second time derivative of the velocity is proportional to the effective social force, instead of the first time derivative as in Vicsek. Although it is true that in the thermodynamic limit the large-scale behavior is described by an over-damped theory Cavagna et al. (2019a), namely the Toner-Tu Toner and Tu (1998) hydrodynamic theory (which is a coarse-grained version of the Vicsek model), inertial effects can be observed in finite systems Cavagna et al. (2015). This is relevant in the description of observations of biological flocks, which at sizes of a few thousands of individuals are large but still far enough from the thermodynamic limit that finite-size effects are important. In the language of the renormalization group (RG), a crossover phenomenon arises in the RG flow such that at intermediate sizes the global properties will be described by an inertial fixed point with an unstable direction, instead of the stable over-damped fixed point that rules in the thermodynamic limit Cavagna et al. (2019b).
@@
In this work we consider extending the ISM to include position-dependent forces. In the original ISM the interaction is between velocities (actually, velocity directions), such that particles tend to align the velocities with each other, just as in the Vicsek model. Positions enter the picture only indirectly, through the definition of the interaction network (which can be metric or topological Kumar and De (2021)). There are several reasons why positional forces are desirable. To treat finite systems with metric interactions it is necessary to introduce some kind of confinement, because otherwise small velocity fluctuations eventually lead to particles becoming isolated from the flock and thus “evaporating” the system to infinite dilution. This can be avoided using reflecting boundary conditions, external fields, or inter-particle interactions providing cohesion. Reflecting walls for confinement Chepizhko et al. (2021) or partial confinement Tu et al. (1998); Morin and Bartolo (2018); Moreno et al. (2020), and cohesion through attractive forces Chaté et al. (2008b) have been considered for the Vicsek case. Position-dependent inter-particle interactions may also be included for other purposes, e.g. to add an excluded volume potential that will limit the local density Tu et al. (1998). A position-dependent external field can also be interesting to study perturbation and response: a field coupled to the velocities is straightforward to add (as it has been done for Vicsek Kyriakopoulos et al. (2016)), but experimentally it may be easier to impose a perturbation with a position-dependent field, for example by implementing a moving artificial marker to perturb an insect swarm Attanasi et al. (2014c).
@@ II Inertial Spin model with external position-dependent forces
Originally, the ISM was proposed Attanasi et al. (2014a); Cavagna et al. (2014) based on the experimental observation of second sound (i.e. order-parameter waves) in the turning of starling flocks Attanasi et al. (2014a). Second sound is an indication that an inertial mechanism is at work that results in propagation with constant speed (vs. diffusive propagation as it occurs in the overdamped case). The goal is then to formulate a model with the static properties of Vicsek’s model (i.e. capable of spontaneously producing orientational order) but with inertial rather than diffusive dynamics Cavagna et al. (2018). It seems thus reasonable to seek a Hamiltonian formulation, as this will lead naturally to canonical (i.e. inertial) equations of motion. One then needs to identify the correct canonical coordinate / conjugate momentum pair. A key observation is that when a flock changes direction of motion individual birds turn following paths of approximately the same radius, rather than following parallel paths. This suggests that invariance under internal (rather than global) rotations is the relevant symmetry. Indeed, equal-radius turns are not generated by rigid rotations (as are parallel-path turns), but by rotations of the internal orientation of each individual, which moves at approximately constant speed. These considerations lead to recognize the Vicsek interaction as an interaction between orientations, rather than velocities, and to propose a Hamiltonian formalism in which the particle orientation is the canonical coordinate. Its canonical conjugate, the spin, is the generator of internal rotations.
@@ II.1 The original ISM
Let us first introduce the original ISM and then our proposal to include position-dependent forces. For what follows, it is convenient to derive the ISM using the velocity of the $i$-th particle $\mathbf{v}_{i}$ and its canonical conjugate, instead of orientation and spin. We do not consider speed fluctuations, so that a set of hard constraints
$f_{i}(\mathbf{v}_{i})=\mathbf{v}_{i}^{2}-v_{0}^{2}=0$ (1)
@@
$\mathcal{H}=\sum_{i}\frac{w_{i}^{2}}{2\mu}+\mathcal{V}(\{\mathbf{v}_{i}\})+\sum_{i}\lambda_{i}f_{i}(\mathbf{v}_{i}).$ (4)
@@
where we have added the friciton and stochastic terms, with $\langle\xi_{i}^{\mu}(t)\rangle=0$, $\langle\xi_{i}^{\mu}(t)\xi_{j}^{\nu}(t^{\prime})\rangle=2\eta T\delta_{ij}\delta^{\mu\nu}\delta(t-t^{\prime})$. Note that (11b) and the deterministic part of (11c) can be derived either substituting $\mathbf{s}_{i}$ for $\mathbf{w}_{i}$ in (II.1) or directly from the Hamiltonian (10) and the Poisson brackets
@@ II.2 Position-dependent forces
$\begin{split}\mathcal{H}=&{}\sum_{i}\frac{w^{2}}{2\mu}+\mathcal{V}(\{\mathbf{v}_{i}\})-\sum_{i}\mathbf{v}_{i}\cdot\mathbf{F}(\{\mathbf{r}\})\\ &+\sum_{i}\mathbf{w}\cdot\frac{\mathbf{G}(\{\mathbf{r}\})}{m}+\sum_{i}\lambda_{i}(v_{i}^{2}-v_{0}^{2}),\end{split}$ (13)
@@
$\dot{\mathbf{w}}_{i}$ $=-\frac{\partial\mathcal{H}}{\partial\mathbf{v}_{i}}=-\frac{\partial\mathcal{V}}{\partial\mathbf{v}_{i}}+\mathbf{F}_{i}-2\lambda_{i}\mathbf{v}_{i},$ (14b)
where again to eliminate the $\lambda_{i}$ one uses the fact that first and second time derivatives of the constraints must vanish together with the equations of motion to find
@@
$\quad-\frac{1}{X}\mathbf{G}_{i}\times(\mathbf{v}_{i}\times\mathbf{s}_{i}),$ (17b)
with $\chi=v_{0}^{2}\mu$, $X=v_{0}^{2}m$, which reduce to (II.1) (apart from noise) when $\mathbf{F}_{i}=\mathbf{G}_{i}=0$. Eqs. (II.2) can be further simplified, because not every form of $\mathbf{G}_{i}$ will lead to observable changes in particle trajectories. The situation is similar to the case of the Hamiltonian formulation of electromagnetic forces, where a gauge transformation alters the vector potential but not the trajectories. This can be seen by redefining the momentum and spin. Eqs. (14a) and (17a) suggest to define $\mathbf{z}_{i}=\mathbf{w}_{i}+(\mu/m)\mathbf{G}_{i}$, and $\tilde{\mathbf{s}}_{i}=\mathbf{v}_{i}\times\mathbf{z}_{i}$. Renaming $\tilde{\mathbf{s}}_{i}\longrightarrow\mathbf{s}_{i}$ and adding the stochastic terms the equations of motion finally read
@@
The final equations of motion cannot be written from Poisson brackets, because $\mathbf{r}_{i}$ is not a canonical coordinate, and in consequence $\dot{\mathbf{G}}_{i}=(\mathbf{v}_{i}\cdot\nabla)\mathbf{G}_{i}$ does not follow naturally from $\{\mathbf{G},\mathcal{H}\}$. However it is possible to write a slightly different Hamiltonian that will yield (II.2) from $\dot{\mathbf{v}}_{i}=\{\mathbf{v}_{i},\mathcal{H}\}$, $\dot{\mathbf{s}}_{i}=\{\mathbf{s}_{i},\mathcal{H}\}$, provided $\mathbf{G}_{i}$ can be written as a gradient, $\mathbf{G}_{i}=\nabla\Gamma_{i}$:
@@ II.3 Overdamped limit and dimensionless quantities
It is possible to eliminate $\mathbf{s}_{i}$ and write a single, second-order, equation of motion for the velocity. This can be more convenient for numerical integration, as then one can apply one of the well-known discretizations used in molecular or stochastic dynamics (see e.g. Cavagna et al., 2016)
$\chi\frac{\ddot{\mathbf{v}}_{i}}{v_{0}}+\chi\frac{(\dot{\mathbf{v}}_{i})^{2}}{v_{0}^{2}}\frac{\mathbf{v}_{i}}{v_{0}}+\eta\frac{\dot{\mathbf{v}}_{i}}{v_{0}}=\\ J\sum_{j}n_{ij}\frac{\mathbf{v}_{j}^{\perp}}{v_{0}}+v_{0}\mathbf{F}^{\perp}+\frac{\chi}{X}\left(\frac{\mathbf{v}_{i}}{v_{0}}\cdot\nabla\right)v_{0}^{2}\mathbf{G}_{i}^{\perp}+\bm{\xi}$ (20)
@@
$\frac{\chi}{a^{2}}\ddot{\hat{\bm{\psi}}}_{i}+\frac{\chi}{a^{2}}(\dot{\hat{\bm{\psi}}}_{i})^{2}\hat{\bm{\psi}}_{i}+\frac{\eta}{a}\dot{\hat{\bm{\psi}}}_{i}=\\ J\sum_{j}n_{ij}\hat{\bm{\psi}}_{j}^{\perp}+v_{0}\mathbf{F}^{\perp}+\frac{\chi}{X}\left(\hat{\bm{\psi}}_{i}\cdot\nabla\right)v_{0}^{2}\mathbf{G}_{i}^{\perp}+\sqrt{\frac{2T\eta}{a}}f(\hat{t}),$ (21)
where $\langle f(\hat{t})f({\hat{t}}^{\prime})\rangle=\delta(\hat{t}-\hat{t}^{\prime})$. Choosing $a\propto\eta$ one arrives at an equation that will become first order in the limit $\chi/\eta^{2}\to 0$. It is convenient that $a$ have units of time so that $\hat{t}$ is non-dimensional. We choose $a=\eta/J$ , and write the equations of motion as
@@
$\quad+\sqrt{\frac{2T}{J}}f(\hat{t}),$ (22b)
@@
$\quad+\sqrt{\frac{2T}{J}}f(\hat{t}).$ (24b)
In what follows we shall consider the case of a single particle and a harmonic restoring force $\hat{\mathbf{F}}=-k\hat{\mathbf{x}}$. For this force, the overdamped equation (but not the full ISM) reduces to a form that looks like the harmonic oscillator,
$\ddot{\hat{\mathbf{x}}}=-\hat{v}_{0}k\hat{\mathbf{x}}^{\perp}+v_{0}\sqrt{\frac{2T}{J}}f(\hat{t}),$ (25)
@@
$\eta\ddot{\mathbf{x}}=-kJv_{0}\mathbf{x}+kJv_{0}(\bm{\psi}\cdot\mathbf{x})\bm{\psi}+\sqrt{2\eta T}f(t).$ (26)

### v3

@@ Abstract
We propose an extension to the inertial spin model (ISM) of flocking and swarming. The model has been introduced to explain certain dynamic features of swarming (second sound, a lower than expected dynamic critical exponent) while preserving the mechanism for onset of order provided by the Vicsek model. The ISM has only been formulated with an imitation (“ferromagnetic”) interaction between velocities. Here we show how to add position-dependent forces in the model, which allows to consider effects such as cohesion, excluded volume, confinement and perturbation with external position-dependent field, and thus study this model without periodic boundary conditions. We study numerically a single particle with an harmonic confining field and compare it to a Brownian harmonic oscillator and to a harmonically confined active Browinian particle, finding qualitatively different behavior in the three cases.
@@ I Introduction
Although the Vicsek model is successful in explaining many aspects of flocking at the static, or stationary, level, it is not suitable to interpret certain dynamical features related to the presence of inertial effects. The finding of wave-like propagation of direction information during turns in starling flocks Attanasi et al. (2014a) led to the proposal of the inertial spin model (ISM) Cavagna et al. (2014), which we consider here. The model is described in detail in Sec. II.1, but it is essentially the Vicsek model endowed with a Hamiltonian-like (second order) dynamics, so that second time derivative of the velocity is proportional to the effective social force, instead of the first time derivative as in Vicsek. Although it is true that in the thermodynamic limit the large-scale behavior is described by an over-damped theory Cavagna et al. (2019a), namely the Toner-Tu Toner and Tu (1998) hydrodynamic theory (which is a coarse-grained version of the Vicsek model), inertial effects can be observed in finite systems Cavagna et al. (2015). This is relevant in the description of observations of biological flocks, which at sizes of a few thousands of individuals are large but still far enough from the thermodynamic limit that finite-size effects are important. In the language of the renormalization group (RG), a crossover phenomenon arises in the RG flow such that at intermediate sizes the global properties will be described by an inertial fixed point with an unstable direction, instead of the stable over-damped fixed point that rules in the thermodynamic limit Cavagna et al. (2019b).
@@
In this work we consider extending the ISM to include position-dependent forces. In the original ISM the interaction is between velocities (actually, velocity directions), such that particles tend to align the velocities with each other, just as in the Vicsek model. Positions enter the picture only indirectly, through the definition of the interaction network (which can be metric or topological Kumar and De (2021)). There are several reasons why positional forces are desirable. To treat finite systems with metric interactions it is necessary to introduce some kind of confinement, because otherwise small velocity fluctuations eventually lead to particles becoming isolated from the flock and thus “evaporating” the system to infinite dilution if the boundary conditions are open. This can be avoided using periodic boundary conditions, reflecting boundaries, external fields, or inter-particle interactions providing cohesion. Reflecting walls for confinement Chepizhko et al. (2021) or partial confinement Tu et al. (1998); Morin and Bartolo (2018); Moreno et al. (2020), and cohesion through attractive forces Chaté et al. (2008b) have been considered for the Vicsek case. Position-dependent inter-particle interactions may also be included for other purposes, e.g. to add an excluded volume potential that will limit the local density Tu et al. (1998). Furthermore, an effective spatial confinement mechanism, mediated by external landmarks, is crucial for natural swarms to maintain global cohesion. Incorporating this element into our theoretical model will thus offer a more comprehensive understanding of the biological system. This will help clarify the system’s bulk properties in the presence of confinement as well as the significance of boundaries in the swarm formation. A position-dependent external field can also be interesting to study perturbation and response: a field coupled to the velocities is straightforward to add (as it has been done for Vicsek Kyriakopoulos et al. (2016)), but experimentally it may be easier to impose a perturbation with a position-dependent field, for example by implementing a moving artificial marker to perturb an insect swarm Attanasi et al. (2014c).
@@ II Inertial Spin model with external position-dependent forces
Originally, the ISM was proposed Attanasi et al. (2014a); Cavagna et al. (2014) based on the experimental observation of second sound (i.e. order-parameter waves) in the turning of starling flocks Attanasi et al. (2014a). Second sound is an indication that an inertial mechanism is at work that results in propagation with constant speed (vs. diffusive propagation as it occurs in the overdamped case). The goal is then to formulate a model with the static properties of Vicsek’s model (i.e. capable of spontaneously producing orientational order) but with inertial rather than diffusive dynamics Cavagna et al. (2018). It seems thus reasonable to seek a Hamiltonian formulation, as this will lead naturally to canonical (i.e. inertial) equations of motion. One then needs to identify the correct canonical coordinate / conjugate momentum pair. A key observation is that when a flock changes direction of motion individual birds turn following paths of approximately the same radius, rather than following parallel paths (see Fig. 1). This suggests that invariance under internal (rather than global) rotations is the relevant symmetry. Indeed, equal-radius turns are not generated by rigid rotations (as are parallel-path turns), but by rotations of the internal orientation of each individual, which moves at approximately constant speed. These considerations lead to recognize the Vicsek interaction as an interaction between orientations, rather than velocities, and to propose a Hamiltonian formalism in which the particle orientation is the canonical coordinate. Its canonical conjugate, the spin, is the generator of internal rotations.
Figure 1: Parallel path vs. equal radius turns. Two 3-particle flocks are shown making a clockwise 90 degree turn (light red circles represent the initial positions of the flock, light blue circles the final positions). In the parallel path case (left), the result is a rigid body rotation around a common center; this rotation is generated by usual angular momentum. Instead, in the equal radius case (right), which is required if all particles move at the same speed, the final configuration does not result from a rigid body rotation: note that the particle at the front of the flock is $c$ before the turn and $a$ afterwards. This transformation is generated by the spin, which rotates particles’ velocities.
@@ II.1 The original ISM
Let us first introduce the original ISM and then our proposal to include position-dependent forces. For what follows, it is convenient to derive the ISM using the velocity of the $i$-th particle $\mathbf{v}_{i}$ and its canonical conjugate, instead of following the original presentation Attanasi et al. (2014a); Cavagna et al. (2014, 2018) employing orientation and spin. We do not consider speed fluctuations, so that a set of hard constraints
$g_{i}(\mathbf{v}_{i})=\mathbf{v}_{i}^{2}-v_{0}^{2}=0$ (1)
@@
$\mathcal{H}=\sum_{i}\frac{w_{i}^{2}}{2\mu}+\mathcal{V}(\{\mathbf{v}_{i}\})+\sum_{i}\lambda_{i}g_{i}(\mathbf{v}_{i}).$ (4)
@@
where we have added the friciton and stochastic terms, with $\langle\xi_{i}^{\mu}(t)\rangle=0$, $\langle\xi_{i}^{\mu}(t)\xi_{j}^{\nu}(t^{\prime})\rangle=2\eta T\delta_{ij}\delta^{\mu\nu}\delta(t-t^{\prime})$. Note that (11b) and the deterministic part of (11c) can be derived either substituting $\mathbf{s}_{i}$ for $\mathbf{w}_{i}$ in (6) or directly from the Hamiltonian (10) and the Poisson brackets
@@
Let us stress again that the Hamiltonian structure of the ISM involves $\mathbf{v}_{i}$ and $\mathbf{w}_{i}$ (or $\mathbf{s}_{i}$), and not the positions. For example, in 2-$d$ one can show that the velocity orientations interact like in an inertial Kuramoto model Markou (2021). Adding an external force that acts directly on the velocities is thus as straightforward as adding an external field in a Kuramoto model: one can add a term $\mathbf{h}\cdot\sum_{i}\mathbf{v}_{i}$ in the Hamiltonian, and this fits in perfectly within the Hamiltonian equations, because $\mathbf{v}_{i}$ is a canonical variable. On the other hand, $\mathbf{r}_{i}$ is not a canonical variable, but the integral of $\mathbf{v}_{i}$, i.e. something analogous to the integral of the Kuramoto orientations, which is a non-trivial object to treat within the Hamiltonian formalism. For this reason, adding a force that depends on $\mathbf{r}_{i}$ is not straightforward. This is what we discuss next.
@@ II.2 Position-dependent forces
$\begin{split}\mathcal{H}=&{}\sum_{i}\frac{w^{2}}{2\mu}+\mathcal{V}(\{\mathbf{v}_{i}\})-\sum_{i}\mathbf{v}_{i}\cdot\mathbf{F}(\{\mathbf{r}\})\\ &+\sum_{i}\mathbf{w}_{i}\cdot\frac{\mathbf{G}(\{\mathbf{r}\})}{m}+\sum_{i}\lambda_{i}(v_{i}^{2}-v_{0}^{2}),\end{split}$ (13)
@@
$\dot{\mathbf{w}}_{i}$ $=-\frac{\partial\mathcal{H}}{\partial\mathbf{v}_{i}}=-\frac{\partial\mathcal{V}}{\partial\mathbf{v}_{i}}+\mathbf{F}_{i}-2\lambda_{i}\mathbf{v}_{i}.$ (14b)
So far our choice of the new terms in the Hamiltonian has been motivated by how the equation of motion will look, i.e. whether the force will act directly on the acceleration $\dot{\mathbf{v}}_{i}$, or indirectly through $\dot{\mathbf{w}}_{i}$. However, looking directly at the Hamiltonian one may wonder whether more general terms, coupling non-linearly to $\mathbf{v}_{i}$ or $\mathbf{w}_{i}$, would be better choices. The coupling to $\mathbf{v}_{i}$ is standard in Hamiltonian systems, since $\mathbf{v}_{i}$ plays the equivalent of a position, and can certainly be made non-linear. However, this is clearly not what we want here, because that would lead to forces dependent on both $\mathbf{v}_{i}$ and $\mathbf{r}_{i}$, and we wish to separate the interactions between velocities (the term $\mathcal{V}(\{\mathbf{v}_{i}\})$) and those dependent on positions. As for the term proportional to $\mathbf{w}_{i}$, it turns out that it would be problematic to choose it non-linear in $\mathbf{w}_{i}$. One needs that the terms involving $\mathbf{w}_{i}$ (the “kinetic energy”) be convex, so that the Hamiltonian can be related to a corresponding Lagrangian via a Legendre transformation. A non-convex Lagrangian leads to multi-valued Hamiltonian, and ill-defined dynamics Chi and He (2014). Although non-convex Lagrangians have been discussed for quantum systems and to study spontaneous breaking of time-translation invariance Henneaux et al. (1987); Shapere and Wilczek (2012); Chi and He (2014), we here we’d rather stick to the classical Hamiltonian formalism. Non-linear couplings can be introduced at the price of restricting the parameters (e.g. the coefficients of a polynomial in $\mathbf{w}_{i}$) so that convexity is preserved (as for example in Casiulis et al. (2020), where a Hamiltonian with terms both linear and quadratic in the velocities is considered), we avoid this because it could lead to imposing restrictions on the form of $\mathbf{G}(\{\mathbf{r}_{i}\})$.
We thus restrict ourselves to linear couplings to $\mathbf{v}_{i}$ and $\mathbf{w}_{i}$, and proceed to eliminate the $\lambda_{i}$ from the equations of motion using the fact that first and second time derivatives of the constraints must vanish to find
@@
$-\frac{1}{X}\mathbf{G}_{i}\times(\mathbf{v}_{i}\times\mathbf{s}_{i}),$ (17b)
with $\chi=v_{0}^{2}\mu$, $X=v_{0}^{2}m$, which reduce to (11) (apart from noise) when $\mathbf{F}_{i}=\mathbf{G}_{i}=0$. Eqs. (17) can be further simplified, because not every form of $\mathbf{G}_{i}$ will lead to observable changes in particle trajectories. The situation is similar to the case of the Hamiltonian formulation of electromagnetic forces, where a gauge transformation alters the vector potential but not the trajectories. This can be seen by redefining the momentum and spin: Eqs. (14a) and (17a) suggest to define $\mathbf{z}_{i}=\mathbf{w}_{i}+(\mu/m)\mathbf{G}_{i}$, and $\tilde{\mathbf{s}}_{i}=\mathbf{v}_{i}\times\mathbf{z}_{i}$. Renaming $\tilde{\mathbf{s}}_{i}\longrightarrow\mathbf{s}_{i}$ and adding the stochastic terms the equations of motion finally read
@@
The final equations of motion cannot be written from Poisson brackets, because $\mathbf{r}_{i}$ is not a canonical coordinate, and in consequence $\dot{\mathbf{G}}_{i}=(\mathbf{v}_{i}\cdot\nabla)\mathbf{G}_{i}$ does not follow naturally from $\{\mathbf{G},\mathcal{H}\}$. However it is possible to write a slightly different Hamiltonian that will yield (18) from $\dot{\mathbf{v}}_{i}=\{\mathbf{v}_{i},\mathcal{H}\}$, $\dot{\mathbf{s}}_{i}=\{\mathbf{s}_{i},\mathcal{H}\}$, provided $\mathbf{G}_{i}$ can be written as a gradient, $\mathbf{G}_{i}=\nabla\Gamma_{i}$:
@@ II.3 Overdamped limit and dimensionless quantities
It is possible to eliminate $\mathbf{s}_{i}$ and write a single, second-order, equation of motion for the velocity. This can be more convenient for numerical integration, as then one can apply one of the well-known discretizations used in molecular or stochastic dynamics (see e.g. Cavagna et al., 2016):
$\chi\frac{\ddot{\mathbf{v}}_{i}}{v_{0}}+\chi\frac{(\dot{\mathbf{v}}_{i})^{2}}{v_{0}^{2}}\frac{\mathbf{v}_{i}}{v_{0}}+\eta\frac{\dot{\mathbf{v}}_{i}}{v_{0}}=\\ J\sum_{j}n_{ij}\frac{\mathbf{v}_{j}^{\perp}}{v_{0}}+v_{0}\mathbf{F}^{\perp}+\frac{\chi}{X}\left(\frac{\mathbf{v}_{i}}{v_{0}}\cdot\nabla\right)v_{0}^{2}\mathbf{G}_{i}^{\perp}+\bm{\xi}^{\perp}$ (20)
@@
$\frac{\chi}{a^{2}}\ddot{\hat{\bm{\psi}}}_{i}+\frac{\chi}{a^{2}}(\dot{\hat{\bm{\psi}}}_{i})^{2}\hat{\bm{\psi}}_{i}+\frac{\eta}{a}\dot{\hat{\bm{\psi}}}_{i}=\\ J\sum_{j}n_{ij}\hat{\bm{\psi}}_{j}^{\perp}+v_{0}\mathbf{F}^{\perp}+\frac{\chi}{X}\left(\hat{\bm{\psi}}_{i}\cdot\nabla\right)v_{0}^{2}\mathbf{G}_{i}^{\perp}+\sqrt{\frac{2T\eta}{a}}\bm{\xi}^{\perp}(\hat{t}),$ (21)
where we have redefined $\bm{\xi}$ so that $\langle\bm{\xi}_{i}(\hat{t})\bm{\xi}_{j}({\hat{t}}^{\prime})\rangle=\delta(\hat{t}-\hat{t}^{\prime})\delta_{ij}$. Choosing $a\propto\eta$ one arrives at an equation that will become first order in the limit $\chi/\eta^{2}\to 0$. It is convenient that $a$ have units of time so that $\hat{t}$ is non-dimensional. We choose $a=\eta/J$, and write the equations of motion as
@@
$\quad+\sqrt{\frac{2T}{J}}\bm{\xi}^{\perp}(\hat{t}),$ (22b)
@@
$\quad+\sqrt{\frac{2T}{J}}\bm{\xi}^{\perp}(\hat{t}).$ (24b)
In what follows we shall consider the case of a single particle and a harmonic restoring force $\hat{\mathbf{F}}=-k_{0}\hat{\mathbf{x}}$. For this force, the overdamped equation (but not the full ISM) reduces to a form that looks like the harmonic oscillator,
$\ddot{\hat{\mathbf{x}}}=-\hat{v}_{0}k_{0}\hat{\mathbf{x}}^{\perp}+v_{0}\sqrt{\frac{2T}{J}}\bm{\xi}^{\perp}(\hat{t}),$ (25)
@@
$\eta\ddot{\mathbf{x}}=-k_{0}Jv_{0}\mathbf{x}+k_{0}Jv_{0}(\bm{\psi}\cdot\mathbf{x})\bm{\psi}+\sqrt{2\eta T}\bm{\xi}^{\perp}(t).$ (26)